OpenStudy (anonymous):

Express in a table the cost of the cab ride given the number of miles provided. Number of miles:0,1,2,3,4,5 Total cost (dollars): ???? The price is $2.00 a mile and a flat fee of $10.00. Please help!

OpenStudy (anonymous):

With the flat fee would the cost be $10, $12, $14, $16, $18, $20? I'm horrible at math I'm probably way off...

OpenStudy (brrandyn):

Evaluate what's been presented to you. Price = 2.00 per mile, right? Let "x" be the number of miles. The flat rate of 10.00 is added as a constant. The following equation can be modeled (for your given scenario) \[2x+10\] Replace "x" with the number of miles and multiply/add accordingly.

OpenStudy (brrandyn):

Here's an example. Calculate the cost of a cab ride for 5 miles.\[2(5) + 10\]\[2*5 + 10\]\[10 + 10 = $20.00\]Add the constant, otherwise, the flat rate of 10.00 \[$20.00 + $10.00 = $30.00\]

OpenStudy (anonymous):

Oooohhh okay thank you i think i get it now. It also wants me to write an equation it says "Write an equation in slope intercept, point-slope, or standard form. Explain why you chose the form you did"

OpenStudy (texaschic101):

y = 2x + 10.....sub in any number for x (the number of miles) to find y (the total cost) lets say, he/she rode 5 miles...sub in 5 for x y = 2(5) + 10 y = 10 + 10 y = 20 so when he/she rode the cab 5 miles, the cost was $20. lets say, he/she rode 2 miles y = 2(2) + 10 y = 4 + 10 y = 14 then it would cost $14.

OpenStudy (brrandyn):

Made a silly mistake by adding the constant twice, haha. Same idea. Using the same equation, we've modeled a slope intercept since the variable is limited to 1 and y (cab cost) is to be solved.

OpenStudy (anonymous):

Oh okay so I was correct the first time?